# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v1

@@ IV Efficient OSD with highly reliable subset reduction
An error bit ${\mathbf{E}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals $T$ or $T+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
An error variable has hard reliability $T$ if BP has a constant output at all iterations; a hard reliability of $T+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity.
The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
@@
In topological codes, such as toric or surface codes, an isolated Pauli error refers to a single-qubit Pauli error (including identity) with no neighboring nonidentity Pauli error within a radius of two qubits. These isolated Pauli errors can be decoded using BP4 with high confidence, contributing to the highly reliable parts in the output of BP.
@@
Equation (10) follows from the parity-check condition (2) as
$\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{S}}\Lambda))\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}^{\top}$
$=$ $\begin{bmatrix}\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}&{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}\end{bmatrix}$
and using the conditions ${\mathbf{z}}^{H}={\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
@@
Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small. To address this, we set the soft reliability threshold $\theta$ very close to one to ensure that highly reliable bits are accurate. Furthermore, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits. If they do not, the algorithm reports a failure.
@@ V Approximate degenerate OSD
In this section, we discuss how degeneracy in quantum codes helps reduce the computation of OSD. Recall that in OSD-$w$, $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates are generated, and the one that is optimal with respect to $\delta$-ADD is chosen as the output. As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset. This new candidate must belong to the same error coset, and thus the bit flipping in the reliable subset corresponds to an operator in the normalizer group. We have the following degenerate error candidate pruning conditions.
Lemma 8.
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
1) When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
2) If each error corresponding to flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
@@
1)
In optimal degenerate decoding ($n$-ADD), different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
@@
Exploiting Lemma 8-1) with $n$-ADD makes OSD4-$w$ more effective. By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{S}}}$ and higher order of OSD can be executed. Consequently, (12) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
For checking whether flipping a bit is nontrivial can be done by checking whether its induced error anticommutes with the permuted logical operators. If an error commutes with all logical operators, it belongs to the stabilizer group. This aligns with the standard procedure in QEC simulations for calculating the logical error rate. We note that the computational complexity of verifying such a bit-flipping event is $O((n-k)k)$. Consequently, checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination. This approach is both effective and relatively efficient for implementing high-order OSD.
However, the complexity of $n$-ADD is not practically feasible. We propose to use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators. For the case of a rotated surface code [24], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$. Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$. Assume that OSD4444-0 generates a list of error candidates $\mathcal{E}$. Then, $4$-ADD outputs:
$\hat{{\mathbf{E}}}=\operatorname*{arg\,max}_{{\mathbf{E}}}\sum_{i=0}^{n-k}\Pr\big{\{}{\mathbf{E}}+{\mathbf{S}}_{i}\big{\}},$ (13)
@@
To achieve optimal execution complexity, 00-ADD is often used. Lemma 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used. However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged. Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Lemma 8-1).
The following provides a sufficient condition to determine whether flipping a reliable bit corresponds to a stabilizer.
Lemma 9.
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output. If the weight of the $j$-th column of ${\mathbf{A}}$ is less than $d-1$, then flipping the $j$-th reliable bit of an error candidate corresponds to multiplying the error candidate by a stabilizer.
Proof.
Suppose that ${\mathbf{E}}=\begin{bmatrix}{\mathbf{F}}\ \Big{|}&{\mathbf{E}}_{1}^{R}&{\mathbf{E}}_{2}^{R}&\cdots&{\mathbf{E}}_{n+k}^{R}\end{bmatrix}$ is a valid error candidate, where ${\mathbf{E}}_{i}\in\{0,1\}$ for $i=1,\dots,n$ and ${\mathbf{F}}\in\{0,1\}^{n-k}$. By (9), flipping the bit ${\mathbf{E}}_{j}^{T}$ corresponds to adding ${\mathbf{E}}$ by the vector $\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$. This vector corresponds to a Pauli operator of weight at most $d-1$ and, therefore, is a stabilizer since the minimum distance of the code is $d$.
By Lemma 8-2) and Corollary 9, we establish the following operational degenerate error candidate pruning condition.
@@
Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm. Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{A}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{A}$ is associated with the reliable bits. If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
@@
We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD. It is important to note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Lemma 9 can be integrated into Algorithm 4.
@@
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
@@
else HRSR returns ${\mathbf{z}}^{\prime}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
     if all the columns of $\tilde{{\mathbf{S}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
@@
     Use OSD4-$w$ on ${\mathbf{z}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{E}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}\right)$.
@@
Remark 11.
We can incorporate the idea of checking the commutation relations with permuted logical operators for column-induced errors of weight at least $d$ into Algorithm 4.
For instance, we first identify columns with weights less than $d-1$ and then verify the commutation relations between the permuted logical operators and the errors induced by the remaining columns.
As will be shown in Table VII, using the strategy outlined in Corollary 10, OSD-00 is applied in the majority of trials when the error rate is low.
@@
MBP4+ADOSD4 this paper based on Algorithm 4
MBP4+OSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 4
MBP4+mOSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 5
AMBP4 [31] No postprocessing. Can act as a predecoder for ADOSD4 and OSD4
BP+mOSD-($w$,$\lambda$) [42] binary BP+ order-w OSD4 with additional flips on $\lambda$ positions
BP-OSD-CS($\lambda$) [43] binary BP-OSD-0 with a greed search on (${\lambda\choose 2}$ positions
BP+GDG [54] binary BP with guided decimal guessing based on a decision tree
@@
VI Simulation results
We simulate the performance of various BP+OSD schemes on LCS codes [59], GHP codes [42, 43], BB codes [55], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [22, 23, 24, 25, 26, 58, 28] (see [64, Table I]).
The decoding performance of a quantum code ${\cal C}({\cal S})$ and its decoder is assessed through the logical error rate (LER). An error $E\in{\cal G}_{n}$ is classified as a logical error if it is a nontrivial logical operator.
The simulation parameters are as follows: The maximum number of BP iterations, $T$, is set to $100$ unless otherwise specified. In most cases, a much smaller value of $T$ suffices for good decoding performance due to the rapid convergence of MBP4[31], which typically completes in just a few iterations for codes with lengths under one thousand. For instance, as shown in Table V, When MBP4 successfully decodes the $[[225,1,15]]$ rotated surface code, it requires an average of only $2.41$ iterations at a depolarizing rate of $\epsilon=3.3\%$. Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B.
For each data point in the plots, at least 100 logical error events are collected.
In Sections VI-D and VI-E, we first demonstrate the performance of the proposed BP+OSD schemes. Then, we analyze the execution time for these schemes in Section VI-G, illustrating the efficiency of ADOSD.
In the simulations of ADOSD4, $\Gamma$ is chosen according to (11) for the purpose of comparison with OSD-2. The soft reliability threshold is set to $\theta=0.999995$. Based on our simulations, selecting $\theta$ within the range of $0.999$ to $0.999995$ results in similar performance.
For comparison, we consider the following decoders: (1) the AMBP4 decoder [31] using a random serial schedule. (2) mOSD-($w$, $\lambda$), a partial order-$w$ OSD approach introduced in [42], which selects $\sum_{j=0}^{w}\binom{\lambda}{j}$ candidates for $\lambda$ less reliable bits based on the reliability metric Max. (3) the BP-OSD-CS($\lambda$) decoder in [43]. (4) BP-GDG algorithm in [54]. These decoders are summarized in Table I.
We observe that both parallel and serial schedules in BP yield similar results for toric codes and surface codes. However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code. We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process. Therefore, we consider the parallel schedule for the following simulations. Similar observations are also made in [54, 65].
@@ VI Simulation results
In this subsection, we demonstrate how to optimize the parameters of our BP+OSD schemes using LCS codes [59]. LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions. Decoding LCS codes presents the most challenging problem among all the quantum codes considered in this paper. These codes cannot be effectively handled by our previous AMBP4 decoder [31], even though AMBP4 outperforms binary BP-OSD decoders [42, 43] on all other codes.
In the following, we consider the MBP4 decoder introduced in [31] as a predecoder for OSD. MBP4 includes a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced. For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [12]. We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
Thus, optimizing the choice of $\alpha$ is crucial. For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance. However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance. This optimization should be performed for each specific code, although the same value of $\alpha$ typically applies across an entire code family.
Figure 2 demonstrates the performance of MBP4+ADOSD4444 on the $[[175,7,7]]$ LCS code with various values of $\alpha$. The performance of MBP4+ADOSD4444 improves significantly and stabilizes as $\alpha$ increases from $1$ to approximately $1.6$.
For reference, we provide a curve $y=ax^{4}$ for some $a$ if errors of weight up to half of the code distance $\lfloor\frac{d-1}{2}\rfloor$ can be corrected. Therefore, MBP4+ADOSD4444 with $\alpha=1.6$ matches this prediction around LER $10^{-7}$. On the other hand, AMBP meets an earlier error floor around LER $10^{-3}$.
Figure 2 also includes a curve labeled MLE (independent X/Z), which is an estimate of the maximum likelihood binary decoding performance using integer programming, as reported in [59]. Since the decoding of X and Z errors is performed independently, this MLE curve represents the optimal performance only under the assumption that X and Z errors are uncorrelated. In [59], the performance of the BP-OSD-CS decoding algorithm from [43] was shown to approximately match the MLE curve. However, it is evident that MBP4+ADOSD4 significantly outperforms this MLE curve.

### v2

@@ IV Highly reliable subset reduction
An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals ${\mathrm{T}}$ or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
An error variable has hard reliability ${\mathrm{T}}$ if BP has a constant output at all iterations; a hard reliability of ${\mathrm{T}}+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity. The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
@@
Equation (15) follows from the parity-check condition (2) with
$\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{H}}\Lambda))\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}^{\top}=$ $\begin{bmatrix}\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\\ {\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\end{bmatrix}$
and the conditions ${\mathbf{s}}^{H}={\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
@@
Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small. To address this, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits. If they do not, the algorithm reports a failure.
@@ V Approximate degenerate OSD
In this section, we discuss how degeneracy in quantum codes helps reduce the computation of OSD. Recall that in OSD-$w$, $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates are generated, and the one that is optimal with respect to $\delta$-ADD is chosen as the output. As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset. This corresponds to adding a logical operator in the normalizer group to the error candidate.
By examining these logical operators, we have the following degenerate error candidate pruning conditions.
Theorem 8.
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in (7), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
  1. 1.
    When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
  2. 2.
    If each logical operator induced by flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
@@
1) In $n$-ADD, different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
@@
Exploiting Theorem 8-1) with $n$-ADD makes OSD4-$w$ more effective. By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed. Consequently, (17) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
Proposition 9.
Checking whether a logical operator induced by flipping a reliable bit of the OSD4-0 output is a stabilizer can be done in $O((n-k)k)$ time.
Proof.
Consider an $[[n,k,d]]$ stabilizer code with a $m\times{2n}$ check matrix ${\mathbf{H}}$ and a $2k\times 2n$ binary logical matrix ${\mathbf{L}}$. During the process of OSD4-0, ${\mathbf{H}}$ leads to $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (7) using column permutations $\pi,\mu$ and row operations $R$, where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
By (13), flipping the $j$-th reliable bit corresponds to adding the vector ${\mathbf{g}}=\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$. The vector ${\mathbf{g}}$ corresponds to a stabilizer if $\mu(\pi({\mathbf{L}}\Lambda)){\mathbf{g}}^{\top}=\mathbf{0}$, and otherwise, it represents a nontrivial logical operator. Since ${\mathbf{A}}_{j}$ is of dimension $n-k$, this computation takes $O((n-k)k)$ time.
The computational complexity of verifying such a bit-flipping event is $O((n-k)k)$, which implies that checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination. This approach is both effective and relatively efficient for implementing high-order OSD.
However, the complexity of $n$-ADD is not practically feasible. We may use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators. For the case of a rotated surface code [25], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$. Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$. Assume that OSD4444-0 generates a list of error candidates $\mathcal{E}$. Then, $4$-ADD outputs:
$\hat{{\mathbf{e}}}=\operatorname*{arg\,max}_{{\mathbf{e}}}\sum_{i=0}^{n-k}\Pr\big{\{}{\mathbf{e}}+{\mathbf{H}}_{i}\big{\}},$ (18)
@@
To achieve optimal execution complexity, 00-ADD is often used. Theorem 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used. However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged. Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Theorem 8-1).
However, Theorem 8-2) still works and we may use Proposition 9 to check this condition.
In the following we provide a weaker sufficient but operationally more efficient condition, for degenerate error candidate pruning.
@@
Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm. Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{{\mathbf{A}}}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{{\mathbf{A}}}$ is associated with the reliable bits. If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
@@
We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD. Note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Proposition 9 can be integrated into Algorithm 4.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
@@
else HRSR returns ${\mathbf{s}}^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
     if all the columns of $\tilde{{\mathbf{H}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
@@
     Use OSD4-$w$ on ${\mathbf{s}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{e}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
@@
VI Simulation results
We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [60], GHP codes [43, 44], BB codes [56], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [23, 24, 25, 26, 27, 59, 29] (see [64, Table I]).
We use the MBP4 decoder introduced in [32] as a predecoder for OSD. MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced. For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13]. We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used. The optimization of $\alpha$ is discussed in Section VI-A.
The maximum number of BP iterations, ${\mathrm{T}}$, is set to $100$ unless otherwise specified. Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B. The soft reliability threshold is set to $\theta=0.999995$ and its selection is discussed in Section VI-C. In ADOSD4 simulations, $\Gamma$ is chosen according to (16) to enable comparison with OSD-2.
In Section VI-D, we compare the proposed BP+OSD schemes with various reliability measures using GHP codes. We then evaluate decoding performance on BB codes and topological codes in Sections VI-E, and VI-F. Finally, Section VI-G analyzes the execution time of these schemes, demonstrating the efficiency of ADOSD.
The decoding performance of a quantum code ${\cal C}({\cal S})$ and a decoder is assessed through the logical error rate (LER). For a code with distance $d$, errors up to $t=\lfloor\frac{d-1}{2}\rfloor$ are correctable, while some errors of weight $t+1$ become uncorrectable. Consequently, at low physical error rates, the performance curve of a decoder follows a scaling behavior of $O(\epsilon^{t+1})$. This implies that in a log-domain plot, the decoder’s performance curve should exhibit an error floor with a slope of $t+1$. A good decoder will have a low error floor. However, if a decoder fails to correct all errors of weight up to $t$, the error floor will have a slope lower than $t+1$. In the worst case, a flatter error floor may appear earlier than expected.
Each data point in the plots is based on at least 100 logical error events in the low error rate regime. At LER $10^{-6}$, this represents an error bar less than $2\times 10^{-7}$, which is roughly the size of the markers for the data points. Thus the error bars are neglected for clarity.
We observe that both parallel and serial schedules in MBP4+ADOSD4 yield similar results for toric codes and surface codes. However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code. We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process. Therefore, we consider the parallel schedule for the following simulations of MBP4+ADOSD4. Similar observations are also made in [55, 65].
@@ VI Simulation results
MBP4+ADOSD4 based on Algorithm 4
MBP4+OSD4-$w$ this paper reliability metric based on Definition 4
MBP4+mOSD4-$w$ reliability metric based on Definition 5
AMBP4 [32] No postprocessing. Can act as
a predecoder for ADOSD4 and OSD4
BP+mOSD($w$,$\lambda$) [43] binary BP+ order-w OSD4
with additional flips on $\lambda$ positions
[…]
a greedy search on (${\lambda\choose 2}$ positions
BP+GDG [55] binary BP with guided decimation
guessing based on a decision tree
@@
For comparison, we consider several decoders listed in Table I. Except for AMBP, we do not implement these decoders ourselves but instead compare our results with those reported in the literature.
In [66], it was shown that a random order schedule improves BP decoding performance. Here, we implement the AMBP decoder with a random order serial schedule, achieving better performance than the results reported in [32]. Note that AMBP incorporates various techniques for performance enhancement, and we do not focus on optimizing its performance in the following simulations.
@@
In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [60] with ${\mathrm{T}}=100$ and $\theta=0.999995$. For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance. However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance. This optimization should be conducted for each code and potentially for each error rate, though a single value of $\alpha$ often applies across an entire code family.
LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions. Decoding LCS codes presents the most challenging problem for AMBP4 among all the quantum codes considered in this paper. We find that the AMBP4 decoder [32] is ineffective for LCS codes, despite its superior performance over binary BP-OSD decoders [43, 44] on all other codes.
Figure 2 compares various decoders on the $[[175,7,7]]$ LCS code, including an MLE (independent X/Z) reference curve from [60]. This MLE curve serves as a benchmark, as a decoder capable of correcting errors of weight up to $3$ should exhibit a performance curve parallel to it at low error rates.
The performance of MBP4+ADOSD4 on the $[[175,7,7]]$ LCS code, with varying values of $\alpha$, exhibits fluctuations at lower values of $\alpha$ but stabilizes as $\alpha$ increases to 1.6. With $\alpha=1.6$, its performance curve indicates that it can correct nearly all weight-3 errors at LER $10^{-7}$.
In [60], the BP-OSD-CS60 decoder was shown to closely match the MLE curve on this code. Since the decoding of $X$ and $Z$ errors is performed independently, the MLE curve represents optimal performance under the assumption that $X$ and $Z$ errors are uncorrelated. As MBP4+ADOSD4 leverages the $X/Z$ correlations, it significantly outperforms both the MLE curve and BP-OSD-CS60 by more than 15 times. In contrast, AMBP4 encounters an error floor for $\epsilon\leq 0.03$.
@@
For improved performance, $\alpha$ can be adaptively chosen for each $\epsilon$, as demonstrated in Section VI-D.